OpenStudy (anonymous):

(9a^2)/3a+b-b^2/3a+b simplify to lowest terms

OpenStudy (anonymous):

3 a+1/(3 a)

OpenStudy (anonymous):

\[\frac{9a^2}{3a+b}-\frac{b^2}{3a+b}\] like that?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

denominators are the same, so the first step is to add in the numerator, keep the denominator

OpenStudy (anonymous):

second step is to factor the numerator it is the difference of two squares third step is to cancel

OpenStudy (anonymous):

9a^2-b^2/3a+b

OpenStudy (anonymous):

that is step one now you need to factor the numerator

OpenStudy (mayankdevnani):

is it???

OpenStudy (mayankdevnani):

plz...sir give me medal fr that question..

OpenStudy (mayankdevnani):

is that question sir u said

OpenStudy (anonymous):

these last steps are what I am struggling with

OpenStudy (mayankdevnani):

sry!!1

OpenStudy (anonymous):

so whatt is 9a^2-b^2 factored

OpenStudy (mayankdevnani):

trying

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

\(x^2-y^2=(x+y)(x-y)\)

OpenStudy (mayankdevnani):

sry!! can't solve

OpenStudy (anonymous):

ok satellite73 whats the final answer here please